OpenStudy (anonymous):

The City Transit Authority plans to hire 8 new bus drivers. From a group of 100 qualified applicants, of whom 65 are men and 35 are women, 8 names are to be selected by lot. Suppose that Mary and John Lewis are among the 100 qualified applicants. What is the probability that both Mary's and John's names will be selected? (Round your answer to four decimal places.)

OpenStudy (anonymous):

men and women make no difference in this problem

OpenStudy (anonymous):

you need john and mary and 6 others out of the 8 chosen, so your answer is \(\frac{\binom{8}{6}}{\binom{100}{8}}\)

OpenStudy (anonymous):

\[\left(\begin{matrix}98 \\ 6\end{matrix}\right)\]/\[\left(\begin{matrix}100 \\ 8\end{matrix}\right)\]